What is the mass of
       (a) 0.2 mole of oxygen atoms?
       (b) 0.5 mole of water molecules?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a) Mole of Oxygen atoms = 0.2 mole
Molar mass of oxygen atoms = 16 g
Mass of oxygen atoms = 16 × 0.2 = 3.2 g
(b) Mole of water molecule = 0.5 mole
Molar mass of water molecules = 2 × 1 + 16 = 18 g
Mass of H2O = 18 × 0.5 = 9 g

Thanks